OpenStudy (anonymous):

What is the product of two and two over three times six and one over six in simplest form?

OpenStudy (anonymous):

\[\frac{2*2}{3*6*\frac{1}{6}}\] Am I interpreting your problem correctly?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

its 2 2/3 x 6 1/6

OpenStudy (anonymous):

\[2\frac{2}{3} * 6\frac{1}{6}\] Convert your mixed numbers into solely fractions and multiply.

OpenStudy (anonymous):

i got 296 /18

OpenStudy (anonymous):

Correct, but it can be simplified.

OpenStudy (anonymous):

by what ?

OpenStudy (anonymous):

Well, both the numerator and denominator are even, so they can definitely be divided by 2.